OpenStudy (anonymous):

simplify: 5 - 3 ---- -- 2a+1 2a answers: a) 2 --- 2a+1 b) 10a-3 ----- 2a+1 c) 16a+3 ------ 2a(2a+1) d) 4a+3 ------ 2a(2a+1) e) 4a-3 ------- 2a(2a+1)

OpenStudy (anonymous):

5/(2a+1) - 3/(2a) =(5*2a - 3*(2a+1)) / ((2a+1)(2a)) =(10a -6a -3) / ((2a+1)(2a)) =(4a-3) / ((2a+1)(2a))

OpenStudy (anonymous):

d right:p